# Smoothing Variances Across Time: Adaptive Stochastic Volatility

### v4

@@ 2.3 Dynamic Shrinkage Processes and their Properties
Thus, assuming $\mu=0$ and $\phi=0.5$ for simplicity, we have the following density function for $v_{t}$, $\lambda_{t}:=exp\{v_{t}/2\}$ and $\kappa_{t}:=\frac{1}{1+exp\{v_{t}\}}$ for the stationary distribution of DSP:
@@
$f(\kappa_{t})=\frac{1}{8(\kappa_{t})(1-\kappa_{t})}\mathrm{sech}^{2}\bigg{(}\frac{1}{4}\bigg{(}\log(\frac{1-\kappa_{t}}{\kappa_{t}})\bigg{)}\bigg{)}=\frac{1}{\sqrt{\kappa_{t}(1-\kappa_{t})}}\frac{1}{2\kappa_{t}(1+\sqrt{\frac{1-\kappa_{t}}{\kappa_{t}}})^{2}}.$
@@
Figure 2 compares the prior distribution of $\lambda_{t}$ and $\kappa_{t}$ for the horseshoe and DSP. The horseshoe prior in Carvalho et al. (2010) assumes $\lambda_{t}\sim C^{+}(0,1)$ and $\kappa_{t}\sim Beta(1/2,1/2)$. The density function of $g(\lambda_{t})$ and $g(\kappa_{t})$ under the horseshoe are:
@@
For $\lambda_{t}$, the density function for DSP has heavier tails and a sharper spike at zero. Specifically, $f(\lambda_{t})>g(\lambda_{t})$ when $\lambda_{t}\in(0,\frac{2-\sqrt{(4-\pi)\pi}}{(\pi-2)})\cup(\frac{2+\sqrt{(4-\pi)\pi}}{(\pi-2)},\infty)$. In terms of $\kappa_{t}$, the DSP prior puts more weight around 0 and 1 than the horseshoe. This is due to the extra U-shaped term in $\kappa_{t}$, $1/(2\kappa_{t}(1+\sqrt{\frac{1-\kappa_{t}}{\kappa_{t}}})^{2})$, pushing the mass of the distribution to near 0 and 1.
Finally, the marginal prior on $\Delta h$ is considered:
@@
Let $|\phi|=0.5$, $\eta_{t-h}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$,$z_{t,h}:=\phi^{h}\eta_{t-h}$, $v_{t}:=\sum_{h=0}^{\infty}z_{t,h}$, $\lambda_{t}:=exp\{v_{t}/2\}$ and $\Delta h_{t}\sim N(0,\lambda_{t})$,
@@
$K_{L}\log(1+\frac{4}{\Delta^{2}h_{t}})<f(\Delta h_{t})<K_{U}\log(1+\frac{2}{\Delta^{2}h_{t}}),$ $|\Delta h_{t}|>0$
where $K_{U}=\frac{1}{2\sqrt{2\pi}}$ and $K_{L}=\frac{1}{8\sqrt{2\pi}}$
@@
The proof is similar to Theorem 1 in Carvalho et al. (2010). We use the fact that $\forall x>0$,
@@
The bounds for the marginal distribution of DSP are similar to the ones from the horseshoe with the only difference being the constant factor $K_{L}$ and $K_{U}$. Under the horseshoe prior, $K_{U}=\frac{1}{\sqrt{2\pi^{3}}}$ and $K_{L}=\frac{1}{2\sqrt{2\pi^{3}}}$. Like horseshoe prior, DSP is also unbounded near the origin, which leads to super-efficiency in a sparse setting as shown in Carvalho et al. (2010).
@@ 2.4 Adaptive Stochastic Volatility via Global-Local Shrinkage
We propose ASV, which uses DSP explored in Section 2.3, to adaptively estimate time-varying volatility in a smooth yet locally adaptive manner. In ASV, the prior models the log-variance of the $k$th-order differences of $h_{t}$: $\Delta^{k}h_{t}|\sigma^{2}_{h,t}\sim N(0,\sigma^{2}_{h,t})$. As previously defined, $v_{t}:=\log(\sigma^{2}_{h,t})$ follows an autoregressive structure described in Equation 3. Unlike RWSV-BL, ASV incorporates both the global parameter $\mu(1-\phi)$ and the local parameter $\phi v_{t-1}+\eta_{t}$. The global parameter promotes overall smoothness, while the local parameter preserves local adaptivity. The resulting model, ASV, has the following hierarchical representation:
$y_{t}=\exp\{h_{t}\}\epsilon_{t}$ $[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$ (4)
@@
The differencing order $k$ plays a central role in controlling the model’s smoothness. ASV can be viewed as a variance-process analogue of smoothing methods for the mean function, such as the Bayesian Trend Filter by Roualdes (2015) and Kowal et al. (2019), Hodrick-Prescott filter by Hodrick and Prescott (1997), $\ell_{1}$-trend filter by Kim et al. (2009) and Tibshirani (2014), and a smoothing spline. Penalizing the $k$th-order difference encourages the underlying signal to be locally well-approximated by a polynomial of degree $k-1$: $k=1$ promotes piecewise constant behavior, $k=2$ encourages piecewise linearity, and higher-order penalties induce smoother trends. In our study, we use k = 1, consistent with traditional volatility models like GARCH and SV, which typically assume stationary dynamics with abrupt clustering. Nevertheless, higher-order differencing may be better suited for applications with more gradually evolving volatility.
In practice, volatility processes may exhibit both local irregularities that are neither smooth nor abrupt. To account for this, we extend ASV by introducing a nugget effect, an additional noise term that absorbs small scale lodl fluctuations. We refer to this model as ASV with nugget:
$h_{t}=h^{*}_{t}+u_{t}$ $[u_{t}|\sigma^{2}_{c}]\sim N(0,\sigma^{2}_{c}),$ (5)
@@
ASV with nugget described in Equation 5 shares the same observation model in Equation 4, but differs in that the log-variance $h_{t}$ is now decomposed into a smooth component $h_{t}^{*}$ and a local noise term $u_{t}$, allowing the model to capture both structured and unstructured variation.
@@ 3 Bayesian Inference via Data Augmentation
We briefly discuss the Bayesian inference procedure for the ASV model described in Equation 4. The observed data are denoted by $\boldsymbol{y}:=(y_{1},\ldots,y_{T})^{\prime}$, the latent log-variance process by $\boldsymbol{h}:=(h_{1},\ldots,h_{T})^{\prime}$ and its associated log-variance process by $\boldsymbol{v}:=(\log(\sigma^{2}_{h,1}),\ldots,\log(\sigma^{2}_{h,T}))^{\prime}$. For the autoregressive parameter $\phi$, a shifted Beta prior: $(\phi+1)/2\sim Beta(1/2,1/2)$ is assumed. For the global location parameter, we assume $\mu\sim Z(1/2,1/2,0,1)$, which implies $\tau=\exp\{\mu(1-\phi)/2\}\sim C^{+}(0,1)$ when $\phi=0$. This recovers the global scale prior used in the horseshoe formulation by Carvalho et al. (2010). Inference proceeds by sampling from the joint posterior distribution of $(\boldsymbol{h},\boldsymbol{v},\mu,\phi)$ conditional on $\boldsymbol{y}$ via Gibbs sampling scheme that iteratively samples from the full conditional distributions.
Sampling from the conditional posterior distribution of $\boldsymbol{h}$ in the SV model is challenging due to its non-linear likelihood. While methods like sequential Monte Carlo by Jacquier et al. (1994) can sample $h_{t}$ from the exact conditional posterior sequentially, this approach is computationally demanding. Instead, we adopt a quasi-likelihood approach based on the transformed observations $\boldsymbol{y^{*}}:=(\log(y^{2}_{1}),\ldots,\log(y^{2}_{T}))^{\prime}$, recasting the model into a linear system with a non-Gaussian error term: $y_{t}^{*}=h_{t}+\log(\epsilon_{t}^{2}),\quad\epsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$.
The asymmetric error term $\log(\epsilon_{t}^{2})$ can be accurately approximated by a finite Gaussian mixture. The idea is originally proposed by Kim et al. (1998) and later refined by Omori et al. (2007), who introduced a 10-component Gaussian mixture approximation:
$\log(\epsilon_{t}^{2})\approx m_{j_{t}}+o_{t}$ $[o_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,w_{j_{t}}^{2}),$
where $j_{t}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Categorical}(\pi^{\text{Omori}})$, with fixed weights $\pi^{\text{Omori}}=(\pi_{1},\ldots,\pi_{10})$. Each latent index $j_{t}\in\{1,\ldots,10\}$ selects a component-specific mean $m_{j_{t}}$ and standard deviation $w_{j_{t}}$ from a pre-specified table (Omori et al., 2007). This approximation transforms the quasi-likelihood into a conditionally Gaussian model, making posterior sampling significantly more efficient. Conditional on $j_{t}$, the full conditional distribution of $\boldsymbol{h}$ becomes a Gaussian likelihood with a structured Gaussian prior:
$y_{t}^{*}\approx h_{t}+m_{j_{t}}+o_{t}$ $[o_{t}|s_{j_{t}}]\stackrel{{\scriptstyle ind}}{{\sim}}N(0,w_{j_{t}}^{2}),$
@@
The same Gaussian mixture approximation and data augmentation are applied to $\log(({\Delta^{k}h_{t}})^{2})$, which shares the same distribution as $\boldsymbol{y}^{*}$. A corresponding latent variable $s_{t}$ is introduced.
The evolution equation of $v_{t}$ in Equation 3 may be interpreted as the SV model with a heavy-tailed noise distribution $\eta_{t}\sim Z(a,b,0,1)$. We leverage the Pólya-Gamma scale mixture representation of the $Z$-distribution (Barndorff-Nielsen et al., 1982), which states:
@@
Combined with the approximation on $\log(\Delta^{k}h_{t}^{2})$, the log-variance evolution becomes conditionally Gaussian:
$\log((\Delta^{k}h_{t})^{2})\approx v_{t}+m_{s_{t}}+r_{t}$ $[r_{t}|w_{s_{t}}]\stackrel{{\scriptstyle ind}}{{\sim}}N(0,w_{s_{t}}^{2}),$
@@
We sample $\boldsymbol{v}$ using the efficient AWOL sampler proposed by Kastner and Frühwirth-Schnatter (2014). Sampling from the Pólya-Gamma distribution is performed by truncating an infinite sum as shown in (Polson et al., 2013). Similarly, $\xi_{\mu}\sim PG(1,0)$ is introduced to represent $\mu\sim Z(1/2,1/2,0,1)$ as $\mu|\xi_{\mu}\sim N(0,\xi_{\mu}^{-1})$.
In addition to the four existing parameters $(\boldsymbol{h},\boldsymbol{v},\mu,\phi)$, four additional parameters are introduced: the mixture indicators $\boldsymbol{j}:=(j_{1},\ldots,j_{T})^{\prime}$ and $\boldsymbol{s}:=(s_{1},\ldots,s_{T})^{\prime}$, and the Pólya-Gamma variables: $\boldsymbol{\xi}:=(\xi_{1}\ldots\xi_{T})^{\prime}$ and $\xi_{\mu}$. This section shows how the data augmentation framework enables efficient posterior sampling of $\boldsymbol{h}$ and $\boldsymbol{v}$. Detailed derivations of the full conditional distributions are discussed in the Appendix B.
For the nugget model described in Equation 5, the same approximation and data augmentation are applied. An additional Gaussian layer $h_{t}|h_{t}^{*},\sigma^{2}_{c}\sim N(h_{t}^{*},\sigma^{2}_{c})$ is introduced and the prior is placed on $\Delta^{k}h^{*}_{t}$ instead of $\Delta^{k}h_{t}$. However, the resulting model remains conditionally Gaussian, and the structure of the Gibbs sampler remains largely unchanged.

### v5

@@ 2.3 Dynamic Shrinkage Processes and their Properties
Thus, assuming $\mu=0$ and $\phi=0.5$ for simplicity, we have the following density functions for $v_{t}$, $\lambda_{t}:=\exp\{v_{t}/2\}$, and $\kappa_{t}:=1/(1+\exp\{v_{t}\})$, corresponding to the stationary distribution of DSP:
@@
$f(\kappa_{t})=\frac{1}{8(\kappa_{t})(1-\kappa_{t})}\mathrm{sech}^{2}\bigg(\frac{1}{4}\log\bigg(\frac{1-\kappa_{t}}{\kappa_{t}}\bigg)\bigg)=\frac{1}{2}\kappa_{t}^{-3/2}(1-\kappa_{t})^{-1/2}\bigg(1+\sqrt{\frac{1-\kappa_{t}}{\kappa_{t}}}\bigg)^{-2}.$
@@
Figure 2 compares the prior distribution of $\lambda_{t}$ and $\kappa_{t}$ under the horseshoe and DSP priors. The horseshoe prior assumes $\lambda_{t}\sim C^{+}(0,1)$ and $\kappa_{t}\sim\text{Beta}(1/2,1/2)$. The corresponding density functions, $g(\lambda_{t})$ and $g(\kappa_{t})$, are given by :
@@
For $\lambda_{t}$, the DSP density has heavier tails and a sharper spike at zero. Specifically, $f(\lambda_{t})>g(\lambda_{t})$ when $\lambda_{t}\in\bigg(0,\frac{2-\sqrt{(4-\pi)\pi}}{(\pi-2)}\bigg)\cup\bigg(\frac{2+\sqrt{(4-\pi)\pi}}{(\pi-2)},\infty\bigg)$. For $\kappa_{t}$, the DSP prior places more weight near 0 and 1 than the horseshoe. This behavior arises from the additional U-shaped term in $\kappa_{t}$, $(2\kappa_{t})^{-1}\bigg(1+\sqrt{\frac{1-\kappa_{t}}{\kappa_{t}}}\bigg)^{-2}$, which pushes the mass of the distribution toward the boundaries.
Finally, the marginal prior on $\{\Delta h_{t}\}_{t=1}^{T}$ is considered:
@@
Let $|\phi|=0.5$, $\eta_{t-\ell}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$, $z_{\ell,t}:=\phi^{\ell}\eta_{t-\ell}$, $v_{t}:=\sum_{\ell=0}^{\infty}z_{\ell,t}$, $\lambda_{t}:=\exp\{v_{t}/2\}$, and $\Delta h_{t}\sim N(0,\lambda_{t}^{2})$. Then
@@
$K_{L}\log\bigg(1+\frac{4}{(\Delta h_{t})^{2}}\bigg)<f(\Delta h_{t})<K_{U}\log\bigg(1+\frac{2}{(\Delta h_{t})^{2}}\bigg),$ $|\Delta h_{t}|>0$
where $K_{U}=1/(2\sqrt{2\pi})$ and $K_{L}=1/(8\sqrt{2\pi})$
@@
The proof is similar to Theorem 1 in horseshoe_theory. We use the fact that $\forall x>0$,
@@
The bounds for the marginal distribution of DSP are similar to the ones from the horseshoe with the only difference being the constant factor $K_{L}$ and $K_{U}$. Under the horseshoe prior, $K_{U}=1/\sqrt{2\pi^{3}}$ and $K_{L}=1/(2\sqrt{2\pi^{3}})$. Like horseshoe prior, DSP is also unbounded near the origin, which leads to super-efficiency in a sparse setting as shown in horseshoe_theory.
@@ 2.4 Adaptive Stochastic Volatility via Global-Local Shrinkage
We propose ASV, which uses DSP explored in Section 2.3, to estimate time-varying volatility. In ASV, the prior models the log-variance of the $k$th-order differences of $h_{t}$, where $\Delta^{k}h_{t}\mid\sigma^{2}_{h,t}\sim N(0,\sigma^{2}_{h,t})$. As previously defined, $v_{t}:=\log(\sigma^{2}_{h,t})$ follows the autoregressive structure described in Equation 3. Unlike RWSV-BL, ASV incorporates both the global parameter, $\mu(1-\phi)$, and the local parameter, $\phi v_{t-1}+\eta_{t}$. The resulting hierarchical representation of ASV is given by:
@@
The differencing order $k$ plays a central role in controlling the model’s smoothness. ASV can be viewed as a variance-process analogue of smoothing methods for the mean function, such as the Bayesian Trend Filter by btf and dsp, the Hodrick-Prescott filter by hp, the $\ell_{1}$ trend filter by l1tf1 and l1tf2, and the smoothing spline by Sspline. Penalizing the $k$th-order difference encourages the underlying signal to be locally well-approximated by a polynomial of degree k1k-1: $k=1$ promotes piecewise constant behavior, $k=2$ encourages piecewise linearity, and higher-order penalties induce smoother trends. In our study, we use $k=1$, consistent with traditional volatility models like GARCH and SV, which typically assume stationary dynamics with abrupt clustering. Nevertheless, higher-order differencing may be better suited for applications with more gradually evolving volatility.
In practice, volatility processes may exhibit local irregularities that are neither smooth nor abrupt. To account for this, we extend ASV by introducing a nugget effect, an additional noise term that absorbs small-scale local fluctuations. We refer to this specification as ASV with nugget:
$h_{t}=h^{*}_{t}+u_{t}$ $[u_{t}|\sigma^{2}_{c}]\sim N(0,\sigma^{2}_{c}),$
@@
It shares the same observation model as its non-nugget counterpart in Equation 4, but differs in that the log-variance $h_{t}$ is decomposed into a smooth component $h_{t}^{*}$, and a local noise term, $u_{t}$, allowing the model to capture both structured and unstructured variation.
@@ 3 Bayesian Inference via Data Augmentation
We briefly outline the Bayesian inference procedure for the ASV model described in Equation 4. The observed data are denoted by $\boldsymbol{y}:=(y_{1},\ldots,y_{T})^{\prime}$, the latent log-variance process by $\boldsymbol{h}:=(h_{1},\ldots,h_{T})^{\prime}$ and its associated log-variance innovation process by $\boldsymbol{v}:=(\log(\sigma^{2}_{h,1}),\ldots,\log(\sigma^{2}_{h,T}))^{\prime}$. For the autoregressive parameter $\phi$, we assume a shifted Beta prior, $(\phi+1)/2\sim\mathrm{Beta}(10,2)$ (dsp). This prior favors positive $\phi$, reflecting the persistent nature of volatility dynamic, while discouraging the unrealistic oscillatory behavior that occurs when $\phi<0.$ For the global parameter, we assume $\mu\sim Z(1/2,1/2,0,1)$, which implies $\tau=\exp\{\mu(1-\phi)/2\}\sim C^{+}(0,1)$ when $\phi=0$. This recovers the same prior on the local parameter in horseshoe_theory. Inference proceeds by sampling from the joint posterior distribution of $(\boldsymbol{h},\boldsymbol{v},\mu,\phi)$ conditional on $\boldsymbol{y}$ via a Gibbs sampling scheme that iteratively samples from the full conditional distributions.
Sampling from the conditional posterior distribution of $\boldsymbol{h}$ in the SV model is challenging due to its nonlinear likelihood. While methods such as sequential Monte Carlo by sequentialMH can sample $h_{t}$ from the exact conditional posterior sequentially, this approach is computationally demanding. Instead, we adopt a quasi-likelihood approach based on the transformed observations $\boldsymbol{y^{*}}:=(\log(y^{2}_{1}),\ldots,\log(y^{2}_{T}))^{\prime}$, recasting the model into a linear system with a non-Gaussian error term: $y_{t}^{*}=h_{t}+\log(\epsilon_{t}^{2})$, where $\epsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$.
The asymmetric error term $\log(\epsilon_{t}^{2})$ is approccimated by a finite Gaussian mixture, originally proposed by kimetal and later refined by omorietal using a 10-component Gaussian mixture:
$\log(\epsilon_{t}^{2})\approx m_{j_{t}}+o_{t}$ $[o_{t}|j_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,w_{j_{t}}^{2}),$
where $j_{t}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Categorical}(\pi^{\mathrm{Omori}})$, with fixed weights $\pi^{\mathrm{Omori}}=(\pi_{1},\ldots,\pi_{10})$. Each latent index $j_{t}\in\{1,\ldots,10\}$ selects a component-specific mean $m_{j_{t}}$ and standard deviation $w_{j_{t}}$ from a pre-specified table (omorietal). This approximation transforms the quasi-likelihood into a conditionally Gaussian model, making posterior sampling significantly more efficient. Conditional on $j_{t}$, the full conditional distribution of $\boldsymbol{h}$ becomes a Gaussian likelihood with a structured Gaussian prior:
$y_{t}^{*}\approx h_{t}+m_{j_{t}}+o_{t}$ $[o_{t}|j_{t}]\stackrel{{\scriptstyle ind}}{{\sim}}N(0,w_{j_{t}}^{2}),$
@@
The same Gaussian mixture approximation and data augmentation are applied to the latent innovation term $\omega_{t}:=\Delta^{k}h_{t}$. Specifically, since $\Delta^{k}h_{t}|v_{t}\sim N(0,\exp(v_{t}))$, we can express
$\omega_{t}=\exp\{v_{t}/2\}a_{t}$ $[a_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1).$
Therefore, $\log(\omega_{t}^{2}):=v_{t}+\log(a_{t}^{2})$, where $\log(a_{t}^{2})$ follows the same asymmetric distribution as $\log(\epsilon_{t}^{2})$ for the process $\{h_{t}\}_{t=1}^{T}$. Consequently, the same 10-component Gaussian mixture approximation can be used, introducing a corresponding latent indicator variable $s_{t}$.
The evolution equation of $v_{t}$ in Equation 3 may be interpreted as the SV model with a heavy-tailed error distribution $\eta_{t}\sim Z(a,b,0,1)$. We leverage the Pólya-Gamma scale mixture representation of the $Z$-distribution (zdist), which states:
@@
Combined with the approximation on $\log(\omega_{t}^{2})=\log((\Delta^{k}h_{t})^{2})$, the log-variance evolution becomes conditionally Gaussian:
$\log(\omega_{t}^{2})\approx v_{t}+m_{s_{t}}+r_{t}$ $[r_{t}|s_{t}]\stackrel{{\scriptstyle ind}}{{\sim}}N(0,w_{s_{t}}^{2}),$
@@
Similarly, $\xi_{\mu}\sim PG(1,0)$ is introduced to represent $\mu\sim Z(1/2,1/2,0,1)$ as $\mu\mid\xi_{\mu}\sim N(0,\xi_{\mu}^{-1})$. Sampling of $\boldsymbol{v}$ follows the AWOL sampler by Kastner_2014, and Pólya-Gamma variables are drawn using a truncated an infinite sum as in (polyagamma).
In addition to the four existing parameters $(\boldsymbol{h},\boldsymbol{v},\mu,\phi)$, four additional parameters are introduced: the mixture indicators $\boldsymbol{j}:=(j_{1},\ldots,j_{T})^{\prime}$ and $\boldsymbol{s}:=(s_{1},\ldots,s_{T})^{\prime}$, and the Pólya-Gamma variables: $\boldsymbol{\xi}:=(\xi_{1}\ldots\xi_{T})^{\prime}$ and $\xi_{\mu}$. Detailed derivations of the full conditional distributions are discussed in the Appendix B. For the nugget model described in Equation 5, the same approximation and data augmentation schemes are applied. An additional Gaussian layer, $h_{t}\mid h_{t}^{*},\sigma^{2}_{c}\sim N(h_{t}^{*},\sigma^{2}_{c})$, is introduced, and the prior is placed on $\Delta^{k}h^{*}_{t}$ instead of $\Delta^{k}h_{t}$. The model remains conditionally Gaussian, and the structure of the Gibbs sampler is largely unchanged.